# Large deviations for the maximum of a reducible two-type branching Brownian motion

### v2

@@ 3.1. Upper bound
and
$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}.$
Set
@@
Then
$f_{2}^{\prime}(u)=\beta-1+\frac{\theta^{2}}{\left[\left(\sigma^{2}-1\right)u+1\right]^{2}}\left(\sigma^{2}-1\right).$
Solving $f_{2}^{\prime}(u_{0})=0$, we have
(39) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\theta-1}{\left|\sigma^{2}-1\right|},$
and
$f_{2}(u_{0})=\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
Then we have several subcases.
@@
(40) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
@@
(41) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
@@
(42) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
@@
(43) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
When $\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$, then $f_{2}^{\prime}(0)\geq 0$ and $f_{2}^{\prime}(1)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ and hence
(44) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\sqrt{2}\theta t\right)\leq\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
@@
$\mathbb{P}\left(M_{t}\geqslant\sqrt{2}\theta t\right)$
$\geqslant$ $\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\right\}.$
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2}\theta t\right)$
$=\beta-\frac{\theta^{2}}{\sigma^{2}}.$
By similar reasoning as (21), (26) and (31), we have
@@
$\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\sum_{u\in{\mathcal{N}}_{s}^{1}(x_{t})}\bar{u}\left(t-s,\sqrt{2\beta\theta^{2}\sigma^{2}}\cdot t-x_{t}\right)\right\}\right]{\rm d}s,$
where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (38) that
@@
This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Choose
$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
for $\varepsilon>0$ small enough. Then we have
$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\sqrt{2}\theta t+\varepsilon t\right){\rm d}s.$
Since
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(M_{(1-\varepsilon)t+s}>\sqrt{2}\theta t+\varepsilon t\right)\geq\beta-\frac{\theta^{2}}{\sigma^{2}},$
then we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\geq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
This proves Theorem 1.2 for the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
@@
$x_{0}=\frac{\sqrt{2}\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t,\quad\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0.$
@@
$\geq\int_{0}^{1}\left(1-e^{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
@@
uniform for $s\in[0,1]$. Furthermore, according to (6),
$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
Thus
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1,$
uniform for $s\in(0,1)$. Then one has
@@
(46) $\geq\beta_{0}+\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
@@
(48) $=\frac{-2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta-\beta+\sigma^{2}}{\sigma^{2}-1}$
(49) $\geq\beta-\frac{\theta^{2}}{\sigma^{2}}=f_{2}(1),$
where the last inequality holds for the case that $\beta<1,\sigma^{2}>1$, and
$\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$
@@
Recall $\xi=\frac{\sigma^{2}-\beta}{\sqrt{2(1-\sigma^{2})(\beta-1)}}.$ The arguments are also exactly the same to those in Section 4. One just to replace $\theta$ by $\frac{\xi\theta}{\sqrt{2}}$. Again, We only give an outline here.
@@ 5.1. Upper bound
$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{\xi^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}.\$
@@
(52) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\xi\theta t\right)\leq f_{3}(0)=1-\frac{\xi^{2}\theta^{2}}{2}.$
If $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$, then $f^{\prime}(1)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$ and hence
(53) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}>\xi\theta t\right)\leq f_{3}(1)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
If $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, then $f_{3}^{\prime}(0)>0$ and $f_{3}^{\prime}(1)<0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ and hence
(54) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln P\left(M_{t}>\xi\theta t\right)\leq f_{3}(u_{0})=\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1.$
@@
$\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right),\,\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)\right\}.$
@@
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\xi\theta t\right)$
$=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
By similar reasoning as (21), (26) and (31), we have
@@
where $I_{t}\subset[0,t]$ is an interval, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (50) that
@@
This proves Theorem 1.3 for the cases that $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, since $1-\frac{\xi^{2}\theta^{2}}{2}\geq\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}$ in this case.
5.2.2. The case of $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(1)$.
This case corresponds to the case that $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$. Choose
$I_{t}=[(1-\varepsilon)t,(1-\varepsilon)t+1],\quad x_{t}=\sqrt{2}\theta t+\varepsilon t,$
for $\varepsilon>0$ small enough. Then we have
@@
$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(\varepsilon t-s,-\varepsilon t\right)}\bigg{|}{\mathcal{N}}_{(1-\varepsilon)t+s}^{1}(x_{t})|\geq 1\right)\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\xi\theta t+\varepsilon t\right){\rm d}s.$
Since
$\lim_{\varepsilon\rightarrow 0}\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{(1-\varepsilon)t+s}>\xi\theta t+\varepsilon t\right)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}},$
then we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
This proves Theorem 1.3 for the case that $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$.
5.2.3. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ for $10<u_{0}<$ 1.
Notice that $f_{3}^{\prime}(u_{0})=0$. This corresponds to the case that $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. Choose
$x_{0}=\frac{\xi\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t,\quad\beta_{0}=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}>0.$
Obviously $x_{0}<\xi\theta$ since $u_{0}>0.$ Then
$\geq\int_{0}^{1}\left(1-e^{-\alpha e^{(\beta_{0}-\varepsilon)t}\bar{u}\left(t-u_{0}t-s,\,\xi\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right){\rm d}s.$
@@
uniform for $s\in[0,1]$. Furthermore, according to (6),
$\frac{1}{t}\log|{\mathcal{N}}_{u_{0}t}^{1}(x_{t})|\rightarrow\beta_{0},\quad a.s.$
Thus
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq e^{(\beta_{0}-\varepsilon)t}\right)=1,$
uniform for $s\in(0,1)$. Then one has
(55) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=f_{3}(u_{0})$
(56) $\geq\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}=f_{3}(1),$
where the last inequality holds for the case that $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. This proves Theorem 1.3 for the above case. We are done.

### v4

@@ 3.1. Upper bound
and by Lemma 2.3, we also have
(38) II $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\mathbb{P}\left(\sigma B_{s}+B_{t}-B_{s}\geqslant\sqrt{2\theta^{2}}\cdot t\right)\mathrm{d}s$
(39) $\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{2\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}\mathrm{d}s$
(40) $=\alpha\int_{0}^{1}e^{f_{2}(u)t}\cdot\frac{\sqrt{t}}{\sqrt{2\pi\left((\sigma^{2}-1)u+1\right)}}\mathrm{d}u,$
where
@@
Then we have several cases.
@@
(42) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
@@
(43) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
@@
(44) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq 1-\theta^{2}.$
@@
(45) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq\beta-\frac{\theta^{2}}{\sigma^{2}}.$
When $\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}\leq\theta^{2}\leq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$, then $f_{2}^{\prime}(0)\geq 0$ and $f_{2}^{\prime}(1)\leq 0$. Thus $\max_{0\leq u\leq 1}f_{2}(u)=f_{2}(u_{0})$ with $f_{2}^{\prime}(u_{0})=0$. Then we have
(46) $u_{0}=\frac{\sqrt{\left|\frac{\sigma^{2}-1}{1-\beta}\right|}\cdot\theta-1}{\sigma^{2}-1}$
and hence
(47) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\sqrt{2}\theta t\right)\leq f_{2}(u_{0})=\frac{2\sqrt{(\beta-1)\left(1-\sigma^{2}\right)}\theta+\beta-\sigma^{2}}{1-\sigma^{2}}.$
@@
(48) $\mathbb{P}\left(M_{t}\geqslant\sqrt{2}\theta t\right)$
(49) $\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right),\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)\right\}.$
@@
(50) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\sqrt{2}\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\sqrt{2}\theta t\right)$
(51) $=\beta-\frac{\theta^{2}}{\sigma^{2}}.$
By similar reasoning as (29) and (32), we have
@@
$\geq\int_{I_{t}}1-\mathbb{E}\left[\exp\left\{-\alpha\left|{\mathcal{N}}_{s}^{1}(x_{t})\right|\bar{u}\left(t-s,\sqrt{2}\theta t-x_{t}\right)\right\}\right]{\rm d}s,$
where $I_{t}\subset[0,t]$ is an interval of length $1$, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (41) that
@@
This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Then with (48) and (50) in hand, Theorem 1.2 for the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}$ follows readily.
@@
$x_{0}=\frac{\sqrt{2}\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t.$
@@
$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(t-u_{0}t-s,\sqrt{2}\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right){\rm d}s.$
@@
uniform for $s\in[0,1]$ and
$\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right)=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}},$
uniform for $s\in[0,1]$. Then one has
@@
(53) $\geq\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}}+\left(1-\frac{\left(\sqrt{2}\theta-x_{0}\right)^{2}}{2\left(1-u_{0}\right)^{2}}\right)\left(1-u_{0}\right)$
@@
(55) $=\frac{-2\sqrt{(1-\beta)\left(\sigma^{2}-1\right)}\theta-\beta+\sigma^{2}}{\sigma^{2}-1}.$
@@
Recall $\xi=\frac{\sigma^{2}-\beta}{\sqrt{2(1-\sigma^{2})(\beta-1)}}.$ The arguments are also exactly the same as those in Section 4. One just replace $\theta$ by $\frac{\xi\theta}{\sqrt{2}}$. We give only an outline here.
@@ 5.1. Upper bound
$\text{II}\leq\alpha\int_{0}^{t}e^{\beta s+(t-s)}\cdot\frac{1}{\sqrt{2\pi\left(\sigma^{2}s+t-s\right)}}\cdot e^{-\frac{\xi^{2}\theta^{2}t^{2}}{2\left(\sigma^{2}s+t-s\right)}}{\rm d}s.\$
@@
(57) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)\leq f_{3}(0)=1-\frac{\xi^{2}\theta^{2}}{2}.$
If $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$, then $f_{3}^{\prime}(1)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$ and hence
(58) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)\leq f_{3}(1)=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
If $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$, then $f_{3}^{\prime}(0)>0$ and $f_{3}^{\prime}(1)<0$. Thus $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ with $f_{3}^{\prime}(u_{0})=0$ yields that
and
(60) $\limsup_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(M_{t}\geq\xi\theta t\right)\leq f_{3}(u_{0})=\frac{\sqrt{2}(1-\beta)-\theta^{2}\left(\sigma^{2}-\beta\right)^{2}-1}{\sqrt{2}\left(\sigma^{2}-1\right)}+1.$
@@
(61) $\mathbb{P}\left(M_{t}\geqslant\xi\theta t\right)\geqslant\max\left\{\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right),\,\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)\right\}.$
@@
(62) $\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{1}:X_{t}^{u}\geqslant\xi\theta t\right)$ $=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{t}\geqslant\xi\theta t\right)$
(63) $=\beta-\frac{\xi^{2}\theta^{2}}{2\sigma^{2}}.$
By similar reasoning as (29) and (32), we have
@@
where $I_{t}\subset[0,t]$ is an interval of length $1$, $x_{t}\in{\mathbb{R}}$ is a deterministic function of $t$. Both $I_{t}$ and $x_{t}$ will be determined later from case to case. Recall from (56) that
@@
$\mathbb{P}_{\sigma^{2},\beta}\left(\bar{M}_{\varepsilon t+s}\geq-\varepsilon t\right)\rightarrow 1,\quad\text{as }t\rightarrow\infty.$
@@
This proves Theorem 1.3 for the cases $\theta^{2}\geq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$.
5.2.2. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(1)$.
This case corresponds to the case $\theta^{2}\leq\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}$. Similarly as before, (61) and (62) yield Theorem 1.3 for this case.
5.2.3. The case of $\max_{0\leq u\leq 1}f_{3}(u)=f_{3}(u_{0})$ for $10<u_{0}<$ 1.
Notice that $f_{3}^{\prime}(u_{0})=0$. This corresponds to the case $\frac{2(\beta-1)}{1-\sigma^{2}}\frac{\sigma^{4}}{\xi^{2}}<\theta^{2}<\frac{2(\beta-1)}{1-\sigma^{2}}\frac{1}{\xi^{2}}$. Choose
$x_{0}=\frac{\xi\theta u_{0}\sigma^{2}}{1-u_{0}+u_{0}\sigma^{2}},\quad x_{t}=x_{0}t.$
Obviously $x_{0}<\xi\theta$, since $u_{0}>0.$ Then
@@
$\geq\int_{0}^{1}\left(1-e^{-\alpha\bar{u}\left(t-u_{0}t-s,\,\xi\theta t-x_{0}t\right)}\right)\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right){\rm d}s.$
@@
uniform for $s\in[0,1]$ and
$\lim_{t\rightarrow\infty}\mathbb{P}\left(|{\mathcal{N}}_{u_{0}t+s}^{1}(x_{t})|\geq 1\right)=\beta u_{0}-\frac{x_{0}^{2}}{2\sigma^{2}u_{0}},$
uniform for $s\in[0,1]$. Then one has
(64) $\liminf_{t\rightarrow\infty}\frac{1}{t}\ln\mathbb{P}\left(\exists\,u\in{\mathcal{N}}_{t}^{2}:X_{t}^{u}\geqslant\xi\theta t\right)=f_{3}(u_{0}).$
This proves Theorem 1.3 for the above case. We are done.
Acknowledgement
We are deeply indebted to Haojun Zhao who pointed out an error in the proof of Theorem 1.1 in the previous version.